OpenStudy (anonymous):

order of operations 8 x2^2-48/8 =

OpenStudy (anonymous):

Do you know PEMDAS?

OpenStudy (anonymous):

yes but i never did this with a fraction in it

OpenStudy (anonymous):

that is the same as dividing

OpenStudy (anonymous):

can u show me in steps

OpenStudy (anonymous):

=8 x 2^2 - 48/8 =8 x 4 - 48/8 =8 x 4 - 6 = ..? can you do the rest?

OpenStudy (anonymous):

26 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks